OpenStudy (mackenzie2013):

What is the value of the 23rd term in the sequence 10, 8, 6, 4, ...? -36 176 -34 -218

OpenStudy (anonymous):

a(n) = a(1) + d(n - 1) This uses arithmetic sequence In this question d = -2, n = 23, a(1) = 10 so a(23) = 10 + -2(22) = 10 - 44 = -34

OpenStudy (mackenzie2013):

Wouldnt you take AN= A1=D(n-1)

OpenStudy (anonymous):

So C is correct

OpenStudy (mackenzie2013):

THank you :)

OpenStudy (anonymous):

np